Solve the equation 1 + 2log (x+1) = log (2x+1) = log (2x+1) +log (5x+8)


1
Expert's answer
2021-06-02T15:58:50-0400

"1+2log(x+1)=log(2x+1)+log(5x+8)."

"log\frac{(2x+1)(5x+8)}{(x+1)^2}=1."

"\frac{(2x+1)(5x+8)}{(x+1)^2}=10."

"10x^2+16x+5x+8=10x^2+20x+10."

"x=2."